If the equilibrium constant for the acid dissociation is KA and pKA (-log Ka) =7.0, what is the ratio of [HPO42-] to [H2PO4−] in this buffer system? Please show your calculation.
Answer #1

The molar ratio is calculated using the Henry Hasselbach equation:

[HPO4-2] / [H2PO4-2] = 10 ^ (pH - pKa) = 10 ^ (7 - 7.2) = 0.63

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.